# Mixture of Parrots [Uncaptioned image][Uncaptioned image][Uncaptioned image]: Experts improve memorization more than reasoning

### v1

@@ Abstract
The Mixture-of-Experts (MoE) architecture enables a significant increase in the total number of model parameters with minimal computational overhead. However, it is not clear what performance tradeoffs, if any, exist between MoEs and standard dense transformers. In this paper, we show that as we increase the number of experts (while fixing the number of active parameters), the memorization performance consistently increases while the reasoning capabilities saturate.
We begin by analyzing the theoretical limitations of MoEs at reasoning. We prove that there exist graph problems that cannot be solved by any number of experts of a certain width; however, the same task can be easily solved by a dense model with a slightly larger width. On the other hand, we find that on memory-intensive tasks, MoEs can effectively leverage a small number of active parameters with a large number of experts to memorize the data. We empirically validate these findings on synthetic graph problems and memory-intensive closed book retrieval tasks. Lastly, we pre-train a series of MoEs and dense transformers and evaluate them on commonly used benchmarks in math and natural language. We find that increasing the number of experts helps solve knowledge-intensive tasks, but fails to yield the same benefits for reasoning tasks.
@@ 2 Related work
Large language models (LLMs) store a considerable amount of knowledge in their parameters (Petroni et al., 2019; Heinzerling & Inui, 2020). They memorize useful knowledge such as facts and commonsense (Zhao et al., 2024). Many works studied how memorization occurs in LLMs by developing tools to locate the knowledge in the model (Meng et al., 2022; Allen-Zhu & Li, 2023; Liu et al., 2024) or by tracking the training dynamics (Tirumala et al., 2022; Speicher et al., 2024). We draw inspiration from Allen-Zhu & Li (2023) and evaluate the memorization of our models by pre-training them on a mixture of datasets that includes Wikipedia, and at test time, evaluate them on world knowledge benchmarks, which are essentially question answering tasks on Wikipedia facts. With respect to theoretical findings, Kim et al. (2023); Mahdavi et al. (2023); Madden et al. (2024) provide upper bounds on the number of parameters needed for dense transformers to perform memorization tasks under various conditions.
@@
In recent years, transformer-based language models have displayed remarkable effectiveness in solving a broad range of reasoning tasks. Specifically, the reasoning capabilities of transformers have been studied in the context of arithmetic problems (Jelassi et al., 2023; Cho et al., 2024; Hou et al., 2024; Zhou et al., 2024; McLeish et al., 2024; Lee et al., 2023), mathematical reasoning (Zhang et al., 2022; Imani et al., 2023; Wei et al., 2022) graph problems (Sanford et al., 2024; Fatemi et al., 2023; Jin et al., 2023; Wang et al., 2024) and code challenges (Shi et al., 2024; Zhu et al., 2024). Recently, state-of-the-art language models were used for solving complex math olympiad problems (DeepMind, 2024; NuminaMath, 2024; OpenAI, 2024). With respect to theoretical findings, various works study the reasoning capabilities of transformers, relating their expressive power to other complexity classes and formal languages (Weiss et al., 2021; Zhou et al., 2023; Strobl et al., 2024). Other works study how chain-of-thought can improve the reasoning capabilities of language models in terms of expressive power and learnability (Abbe et al., 2024; Merrill & Sabharwal, 2023; Malach, 2024). However, the reasoning capabilities of MoE language models compared to their dense counterparts have received comparatively less attention.
@@ 3.2 MoEs require a critical hidden size to solve graph reasoning tasks
Graph connectivity, shortest path, and cycle detection are all graph reasoning tasks which reduce to the length-2 path problem due to (Sanford et al., 2024) and Lemma C.2. We provide a lower-bound on the width required for a sparse transformer to solve the length-2 path problem, and an upper-bound on the width required for a dense transformer to solve the problem. Further, we show a separation between dense and sparse transformers with the same number of parameters: for a sufficiently large amount of experts in the sparse model, it cannot solve the same problem that a dense model can solve with the same amount of total parameters.
@@
The proof follows almost identically from the proof in (Sanford et al., 2024) for the class $\text{Transformer}_{m,H,1}^{N}$. The original proof does not place constraints on the function $\psi$ and is based on a communication-complexity argument. As such we may design $\psi$ so that it first routes and then chooses which expert to apply. We give a complete proof in Appendix C. As such, the result of (Sanford et al., 2024) can also be extended to the class $\text{Transformer}_{m,H,1,K}^{N}$.
@@
The proof relies on an encoding of the inputs where the output values only exceed a certain threshold when $u$ and $v$, the source and destination vertices, have edges with a common vertex. We defer the proof to Appendix C.
@@ 3.3 MoEs use their experts to solve memory-intensive tasks
Specifically, if we choose $K=\sqrt{n/m}$ we get that an MoE architecture can solve the memorization problem with $\tilde{O}(\sqrt{nm})$ parameters. To prove this, we show that for a particular routing function, the number of samples routed to each expert is approximately $n/K$. Then, we show that an expert with $\tilde{O}(n/{mK})$ neurons can memorize a sample of size $O(n/K)$. We present the proof in Appendix C.2.
@@
This bound follows from the fact that there are $2^{n}$ possible labels for any fixed set of $n$ inputs, and at most $2^{cW}$ functions with $W$ parameters and $c$ bit per parameters. The proof is in Appendix C.2.
@@ 4 Synthetic experiments
In the previous section, we proved that there exist graph connectivity problems that cannot be solved by any number of experts of a certain width but the same task can be solved by a dense model with a slightly larger width. Our goal in this section is to verify that our theoretical analysis bears out experimentally when training models from scratch on synthetic data, before moving on to study pre-trained models in Section 5. We mainly focus on two tasks: the shortest path problem (Figure 2), which we use as a synthetic task to represent reasoning problems, and the phone-book task (Figure 3), to measure the memorization ability of our models. Our experiments in this section highlight that adding experts yields greater performance improvements on memorization tasks than reasoning tasks.
@@ 5 Pre-trained Models
In this section, we pre-train dense transformers and MoEs and compare their performance on standard math and natural language benchmarks. We break the downstream tasks into those that require more memorization and those that require more reasoning. The memorization-intensive tasks test for “world knowledge” and consist of benchmarks like TriviaQA (Joshi et al., 2017). We break the reasoning-intensive tasks into two subcategories: one for natural language reasoning tasks like WinoGrande (Sakaguchi et al., 2021) and another for mathematical reasoning tasks like Hendrycks-MATH (Hendrycks et al., 2021). These tasks may be seen as real-world analogs of the stylized phone-book and shortest path tasks studied in Section 4.
@@ 5.1 Setup
We train dense transformers and MoEs using the OLMoE codebase (Muennighoff et al., 2024). We set the number of layers $L=20$ and vary the width $d\in\{256,512,1024,2048,4096\}$ for dense transformers and $d\in\{256,512,1024\}$ for MoEs. Similarly to Muennighoff et al. (2024), we consistently set the intermediate dimension in the FFN/MoE blocks to $d$ (and not $4d$). For MoEs, we vary the number of experts $E\in\{8,16,32,64\}$. For the specific case of width 256, we also train a MoE with 256 experts because its parameter count approximately matches the one of a width-2048 dense model and thus, we can compare the downstream performance of the two models. We use top-2 token-choice routing, without token dropping which is implemented in the dMoE function from the Megablocks package (Gale et al., 2023).
@@
We train two collections of models, one collection on natural language and another one on math. The natural language dataset is a mixture constituted of FineWeb-edu (Penedo et al., 2024), Cosmopedia (Ben Allal et al., 2024), Wikipedia and the training sets of the downstream tasks we evaluate on. The math dataset is a mixture made of Proof-Pile 2 (Azerbayev et al., 2023) and instruction datasets such as OpenMathInstruct (Toshniwal et al., 2024) and MetaMathQA (Yu et al., 2023). Each of the two training mixture approximately totals 65B tokens. A precise description of the training mixtures can be found in Appendix A.
@@
  • Math benchmarks: SVAMP (Patel et al., 2021), GSM8k (Cobbe et al., 2021), GSM-Hard (Gao et al., 2023), Hendrycks-MATH (Hendrycks et al., 2021) and Minerva-MATH (Lewkowycz et al., 2022).
Figure 5: Generalization gap i.e., difference between the training and test accuracies, when the test set is GSM8k (a) and Hendrycks-MATH (b).
@@
In all our experiments, we plot the average accuracy for each of these three categories. We report the corresponding per-task performance in Appendix B.
@@ 5.2 Results
We provide additional evidence that memorization occurs in pre-trained MoEs by considering the generalization gap. In Figure 5 we select 6,319 random problems from the OpenMathInstruct dataset, which is part of the training mixture data. More precisely, we pick 5,000 Hendrycks-MATH like examples and 1,319 GSM8k-like examples to ensure that the number of training examples matches with the corresponding number of examples in GSM8k and Hendrycks-MATH test sets. We then report the generalization gap, which is the gap between the accuracy on training examples and test examples. While both dense transformers and MoEs make a single pass on the OpenMathInstruct dataset, Figure 5 shows that at scales beyond 159M parameters, MoEs suffer from a more significant generalization gap than dense transformers. This is suggestive that MoEs are more prone to memorize training data than dense models.
@@
Finally, we focus on the relationship between validation perplexity and downstream performance in Figure 6. Rather than comparing models by their parameter count, we can compare them based on how well they fit the training distribution as measured by validation perplexity. Even though two models may have the same perplexity, they will have learned different functions. The question is then if we can see any high level patterns in which types of functions a particular model class is more likely to learn. 6(a) shows that at a fixed perplexity, the MoE models outperform the dense models on world knowledge tasks. This suggests that MoEs do have a bias towards learning functions that memorize training data. On the other hand, Figures 6(b) and 6(c) show that MoEs and dense models perform about the same on the reasoning tasks at fixed validation perplexity. We can square this with the results from Figure 1 by noting that at equal total number of parameters an MoE has worse validation perplexity than the corresponding dense model. This suggests that while MoEs do not change the relationship between perplexity and downstream accuracy on reasoning tasks relative to dense models, they may struggle to learn the reasoning parts of the training distribution as well.
Overall, our main findings in Figure 1 and supplementary experiments in Figures 5 and 6 corroborate the hypothesis that MoEs can effectively use more experts to increase their memory capacity, but not necessarily their capability to reason.
@@ 6 Discussion
This paper illuminates the costs and benefits of this reduced dependence on $d$. We show that for some reasoning-intensive tasks increasing the dimension $d$ is inevitable, and scaling the computation with $O(d^{2})$ seems unavoidable. This remains true regardless of the different design choices in the MoE architecture and is backed up empirically. We note that there is increasing interest in developing non-MoE models with sub-quadratic dependence on $d$, using some structural assumptions on the weight layers (Kamalakara et al., 2022; Dao et al., 2021; 2022; Fu et al., 2024), which could provide an alternative.
@@
Limitations and future work.
While we provide substantial experiments on pre-trained models with up to $\leq 2.1$B parameters, we recognize that large scale MoEs like Mixtral (Jiang et al., 2024), DeepSeek-V2 (Dai et al., 2024), and others have orders of magnitude more parameters. We hypothesize that our results would still be meaningful at larger scales due to the strong theoretical underpinning, but it is not guaranteed. Moreover, as suggested above, it would be an interesting direction for future work to propose new architectures with reduced $d$ dependence that can get the best of both worlds and solve reasoning and memorization tasks.
@@ Acknowledgements
Appendix A Details on the pre-training datasets
@@
Appendix B Additional experiments
In all our experiments in Section 5, we report the average accuracy performance obtained by our pre-trained models on respectively world knowledge, commonsense and math benchmarks. Here, we provide the results per task. In Subsection B.1, we display for each task, the downstream performance on a per parameter basis (similar to Figure 1) and in Subsection B.2, we plot for each task, the downstream performance on a per validation perplexity basis (similar to Figure 6).
B.1 Downstream performance on a per parameter basis
@@
B.2 Downstream performance on a per val perplexity basis
@@
Appendix C Proofs
C.1 Reasoning proofs
Definition C.1 (Set-disjointness task).
@@ C.1 Reasoning proofs
Lemma C.2 (Equivalence of set-disjointness and length-2 path).
@@
Lemma C.3 (Communication complexity lower-bound on concatenated outputs).
@@
We complete the first step with Lemma C.3. We complete the second set due to Lemma C.2. It remains to show that a router function also yields the same lower bound. We show that Lemma 39 of Sanford et al. (2024) can be generalized to the case in which $\psi$ is applied according to a routing function. Specifically, consider a top-$1$ routing function $r:\mathbb{R}^{m}\rightarrow[K]$, and $K$ element-wise functions $\psi_{1},\ldots,\psi_{K}:\mathbb{R}^{m}\rightarrow\mathbb{R}$. For shorthand, define:

### v2

@@ Abstract
The Mixture-of-Experts (MoE) architecture enables a significant increase in the total number of model parameters with minimal computational overhead. However, it is not clear what performance tradeoffs, if any, exist between MoEs and standard dense transformers. In this paper, we show that as we increase the number of experts (while fixing the number of active parameters), the memorization performance consistently increases while the reasoning capabilities saturate. We begin by analyzing the theoretical limitations of MoEs at reasoning. We prove that there exist graph problems that cannot be solved by any number of experts of a certain width; however, the same task can be easily solved by a dense model with a slightly larger width. On the other hand, we find that on memory-intensive tasks, MoEs can effectively leverage a small number of active parameters with a large number of experts to memorize the data. We empirically validate these findings on synthetic graph problems and memory-intensive closed book retrieval tasks. Lastly, we pre-train a series of MoEs and dense transformers and evaluate them on commonly used benchmarks in math and natural language. We find that increasing the number of experts helps solve knowledge-intensive tasks, but fails to yield the same benefits for reasoning tasks.
@@ 2 Related work
Large language models (LLMs) store a considerable amount of knowledge in their parameters (Petroni et al., 2019; Heinzerling & Inui, 2020). They memorize useful knowledge such as facts and commonsense (Zhao et al., 2024). Many works studied how memorization occurs in LLMs by developing tools to locate the knowledge in the model (Meng et al., 2022; Allen-Zhu & Li, 2023; Liu et al., 2024) or by tracking the training dynamics (Tirumala et al., 2022; Speicher et al., 2024). We draw inspiration from Allen-Zhu & Li (2023) and evaluate the memorization of our models by pre-training them on a mixture of datasets that includes Wikipedia, and at test time, evaluate them on world knowledge benchmarks, which are essentially question answering tasks on Wikipedia facts. With respect to theoretical findings, Kim et al. (2023); Mahdavi et al. (2023); Madden et al. (2024); Nichani et al. (2024) provide upper bounds on the number of parameters needed for dense transformers to perform memorization tasks under various conditions.
@@
In recent years, transformer-based language models have displayed remarkable effectiveness in solving a broad range of reasoning tasks. Specifically, the reasoning capabilities of transformers have been studied in the context of arithmetic problems (Jelassi et al., 2023; Cho et al., 2024; Hou et al., 2024; Zhou et al., 2024; McLeish et al., 2024; Lee et al., 2023), mathematical reasoning (Zhang et al., 2022; Imani et al., 2023; Wei et al., 2022) graph problems (Sanford et al., 2024; Fatemi et al., 2023; Jin et al., 2023; Wang et al., 2024) and code challenges (Shi et al., 2024; Zhu et al., 2024). Recently, state-of-the-art language models were used for solving complex math olympiad problems (DeepMind, 2024; NuminaMath, 2024; OpenAI, 2024). With respect to theoretical findings, various works study the reasoning capabilities of transformers, relating their expressive power to other complexity classes and formal languages (Weiss et al., 2021; Zhou et al., 2023; Strobl et al., 2024; Chen & Zou, 2024). Other works study how chain-of-thought can improve the reasoning capabilities of language models in terms of expressive power and learnability (Abbe et al., 2024; Merrill & Sabharwal, 2023; Malach, 2024).
@@ 3.2 MoEs require a critical hidden size to solve graph reasoning tasks
Graph connectivity, shortest path, and cycle detection are all graph reasoning tasks which reduce to the length-2 path problem due to (Sanford et al., 2024) and Lemma D.2. We provide a lower-bound on the width required for a sparse transformer to solve the length-2 path problem, and an upper-bound on the width required for a dense transformer to solve the problem. Further, we show a separation between dense and sparse transformers with the same number of parameters: for a sufficiently large amount of experts in the sparse model, it cannot solve the same problem that a dense model can solve with the same amount of total parameters.
@@
The proof follows almost identically from the proof in (Sanford et al., 2024) for the class $\text{Transformer}_{m,H,1}^{N}$. The original proof does not place constraints on the function $\psi$ and is based on a communication-complexity argument. As such we may design $\psi$ so that it first routes and then chooses which expert to apply. We give a complete proof in Appendix D. As such, the result of (Sanford et al., 2024) can also be extended to the class $\text{Transformer}_{m,H,1,K}^{N}$.
@@
The proof relies on an encoding of the inputs where the output values only exceed a certain threshold when $u$ and $v$, the source and destination vertices, have edges with a common vertex. We defer the proof to Appendix D.
@@ 3.3 MoEs use their experts to solve memory-intensive tasks
Specifically, if we choose $K=\sqrt{n/m}$ we get that an MoE architecture can solve the memorization problem with $\tilde{O}(\sqrt{nm})$ parameters. To prove this, we show that for a particular routing function, the number of samples routed to each expert is approximately $n/K$. Then, we show that an expert with $\tilde{O}(n/{mK})$ neurons can memorize a sample of size $O(n/K)$. We present the proof in Appendix D.2.
@@
This bound follows from the fact that there are $2^{n}$ possible labels for any fixed set of $n$ inputs, and at most $2^{cW}$ functions with $W$ parameters and $c$ bit per parameters. The proof is in Appendix D.2.
@@ 4 Synthetic experiments
In the previous section, we proved that there exist graph connectivity problems that cannot be solved by any number of experts of a certain width but the same task can be solved by a dense model with a slightly larger width. Our goal in this section is to verify that our theoretical analysis bears out experimentally when training models from scratch on synthetic data, before moving on to study pre-trained models in Section 5. We mainly focus on two tasks: the shortest path problem (Figure 2), which we use as a synthetic task to represent reasoning problems, and the phone-book task (Figure 3), to measure the recall ability of our models. Our experiments in this section highlight that adding experts yields greater performance improvements on memorization tasks than reasoning tasks.
@@ 5 Pre-trained Models
In this section, we pre-train dense transformers and MoEs and compare their performance on standard math and natural language benchmarks. We break the downstream tasks into those that require more memorization and those that require more reasoning. Here, memorization refers to recall in that we measure the ability of the models to retrieve real-world facts, The memorization-intensive tasks test for “world knowledge” and consist of benchmarks like TriviaQA (Joshi et al., 2017). We break the reasoning-intensive tasks into two subcategories: one for natural language reasoning tasks like WinoGrande (Sakaguchi et al., 2021) and another for mathematical reasoning tasks like Hendrycks-MATH (Hendrycks et al., 2021). These tasks may be seen as real-world analogs of the stylized phone-book and shortest path tasks studied in Section 4.
@@ 5.1 Setup
We train dense transformers and MoEs using the OLMoE codebase (Muennighoff et al., 2024). We set the number of layers $L=20$ and vary the width $d\in\{256,512,1024,2048,4096\}$ for dense transformers and $d\in\{256,512,1024\}$ for MoEs. Similarly to Muennighoff et al. (2024), we consistently set the intermediate dimension in the FFN/MoE blocks to $d$ (and not $4d$). For MoEs, we vary the number of experts $E\in\{8,16,32,64\}$. For the specific case of width 256, we also train a MoE with 256 experts because its parameter count approximately matches the one of a width-2048 dense model and thus, we can compare the downstream performance of the two models. We use top-2 token-choice routing, without token dropping which is implemented in the dMoE function from the Megablocks package (Gale et al., 2023). We leave the study of MoEs trained with other routing mechanisms for future work.
@@
We train two collections of models, one collection on natural language and another one on math. The natural language dataset is a mixture constituted of FineWeb-edu (Penedo et al., 2024), Cosmopedia (Ben Allal et al., 2024), Wikipedia and the training sets of the downstream tasks we evaluate on. The math dataset is a mixture made of Proof-Pile 2 (Azerbayev et al., 2023) and instruction datasets such as OpenMathInstruct (Toshniwal et al., 2024) and MetaMathQA (Yu et al., 2023). Each of the two training mixture approximately totals 65B tokens. A precise description of the training mixtures can be found in Appendix B.
@@
  • Math: SVAMP (Patel et al., 2021), GSM8k (Cobbe et al., 2021), GSM-Hard (Gao et al., 2023), Hendrycks-MATH (Hendrycks et al., 2021), Minerva-MATH (Lewkowycz et al., 2022).
@@
Figure 5: Generalization gap when the test set is GSM8k (a) and Hendrycks-MATH (b).
In all our experiments, we plot the average accuracy for each of these three categories. We report the corresponding per-task performance in Appendix C.
@@ 5.2 Results
We provide additional evidence that memorization occurs in pre-trained MoEs by considering the generalization gap. In Figure 5 we select 6,319 random problems from the OpenMathInstruct dataset, which is part of the training mixture data. More precisely, we pick 5,000 Hendrycks-MATH like examples and 1,319 GSM8k-like examples to ensure that the number of training examples matches with the corresponding number of examples in GSM8k and Hendrycks-MATH test sets. We then report the generalization gap, which is the gap between the accuracy on training examples and test examples. While both dense transformers and MoEs make a single pass on the OpenMathInstruct dataset, Figure 5 shows that at scales beyond 159M parameters, MoEs suffer from a more significant generalization gap than dense transformers. This is suggestive that MoEs are more prone to overfit to the problems they have been pre-trained on than dense models.
@@
Finally, we focus on the relationship between validation perplexity and downstream performance in Figure 6. Rather than comparing models by their parameter count, we can compare them based on how well they fit the training distribution as measured by validation perplexity. Even though two models may have the same perplexity, they will have learned different functions. The question is then if we can see any high level patterns in which types of functions a particular model class is more likely to learn. 6(a) shows that at a fixed perplexity, the MoE models outperform the dense models on world knowledge tasks.
On the other hand, Figures 6(b) and 6(c) show that MoEs and dense models perform about the same on the reasoning tasks at fixed validation perplexity. Note that both dense Transformers and MoEs are trained with the objective of minimizing the perplexity (or loss). However, there could be multiple strategies to achieve the same loss, e.g. by memorizing pieces of data or by improving “reasoning” capabilities. Which strategy the model prefers is determined by the implicit bias of the architecture. In the experiments, we see that for memorization / factual recall, MoEs achieve better accuracy for the same pereplexity value, suggesting that they “prioritize” memorization over other capabilties.
@@ 6 Discussion
This paper illuminates the costs and benefits of this reduced dependence on $d$. We show that for some reasoning-intensive tasks increasing the dimension $d$ is inevitable, and scaling the computation with $O(d^{2})$ seems unavoidable. This remains true regardless of the different design choices in the MoE architecture and is backed up empirically. There is increasing interest in developing non-MoE models with sub-quadratic dependence on $d$, using some structural assumptions on the weight layers (Kamalakara et al., 2022; Dao et al., 2021; 2022; Fu et al., 2024), which could provide an alternative.
@@ Acknowledgements
Appendix A Limitations and future work
While we provide substantial experiments on a wide range of tasks, many of these tasks are not pure memorization nor pure reasoning. We study these two extreme cases to better convey our message but it would be interesting in future work to understand the difference between dense and sparse models on tasks that mix both memorization and recall. Our pre-trained models have up to $\leq 2.1$B parameters, but we recognize that large scale MoEs like Mixtral (Jiang et al., 2024), DeepSeek-V2 (Dai et al., 2024), and others have orders of magnitude more parameters. We hypothesize that our results would still be meaningful at larger scales due to the strong theoretical underpinning, but it is not guaranteed. Moreover, as suggested above, it would be an interesting direction for future work to propose new architectures with reduced $d$ dependence that can get the best of both worlds and solve reasoning and memorization tasks.
Appendix B Details on the pre-training datasets
@@
Appendix C Additional experiments
In all our experiments in Section 5, we report the average accuracy performance obtained by our pre-trained models on respectively world knowledge, commonsense and math benchmarks. Here, we provide the results per task. In Subsection C.1, we display for each task, the downstream performance on a per parameter basis (similar to Figure 1) and in Subsection C.2, we plot for each task, the downstream performance on a per validation perplexity basis (similar to Figure 6).
C.1 Downstream performance on a per parameter basis
@@
C.2 Downstream performance on a per val perplexity basis
@@
Appendix D Proofs
D.1 Reasoning proofs
Definition D.1 (Set-disjointness task).
@@ D.1 Reasoning proofs
Lemma D.2 (Equivalence of set-disjointness and length-2 path).
@@
Lemma D.3 (Communication complexity lower-bound on concatenated outputs).
@@
We complete the first step with Lemma D.3. We complete the second set due to Lemma D.2. It remains to show that a router function also yields the same lower bound. We show that Lemma 39 of Sanford et al. (2024) can be generalized to the case in which $\psi$ is applied according to a routing function. Specifically, consider a top-$1$ routing function $r:\mathbb{R}^{m}\rightarrow[K]$, and $K$ element-wise functions $\psi_{1},\ldots,\psi_{K}:\mathbb{R}^{m}\rightarrow\mathbb{R}$. For shorthand, define: